# Training-Free Constrained Generation with Stable Diffusion Models

### v3

@@ Abstract
Stable diffusion models represent the state-of-the-art in data synthesis across diverse domains and hold transformative potential for applications in science and engineering, e.g., by facilitating the discovery of novel solutions and simulating systems that are computationally intractable to model explicitly. While there is increasing effort to incorporate physics-based constraints into generative models, existing techniques are either limited in their applicability to latent diffusion frameworks or lack the capability to strictly enforce domain-specific constraints. To address this limitation this paper proposes a novel integration of stable diffusion models with constrained optimization frameworks, enabling the generation of outputs satisfying stringent physical and functional requirements. The effectiveness of this approach is demonstrated through material design experiments requiring adherence to precise morphometric properties, challenging inverse design tasks involving the generation of materials inducing specific stress-strain responses, and copyright-constrained content generation tasks.
footnotetext: *Equal and main contribution; 1Polytechnic of Turin; 2University of Genoa; 3University of Virginia. Corresponding authors: Stefano Zampini <stefano.zampini@polito.it>, Jacob K. Christopher <csk4sr@virginia.edu>, Ferdinando Fioretto <fioretto@virginia.edu>
@@ 1 Introduction
However, despite their success in generating coherent content, diffusion models face a critical limitation when applied to domains that require outputs to adhere to strict criteria. In scientific and engineering contexts, generated data must go beyond merely resembling real-world examples; it must rigorously comply with predefined specifications, such as physical laws, first principles, safety standards, or design constraints. When these criteria are not met, the outputs may become unreliable, unsuitable for practical use, or even hazardous, undermining trust in the model’s applicability. Bridging this gap is crucial for realizing the potential of diffusion models in scientific applications.
Recent research has reported varying success in augmenting these models with (often specialized classes of) constraints and providing adherence to desired properties in selected domains [22, 32, 20, 21, 11]. Many of these methods, however, are restricted to simple constraint sets or feasible regions that can be easily approximated, such as a simplex, L2-ball, or polytope. These assumptions break down in scientific and engineering tasks where constraints may be non-linear, non-convex, or even non-differentiable. Moreover, these techniques modify the forward and reverse processes in the original data space, which makes them incompatible with latent diffusion models like Stable Diffusion. Latent-space variants have begun to appear, but they rely on special measurement operators [46, 56, 42] or on learned soft penalties [18, 44], thus limiting their general applicability on the problem of interest to this work
@@ 2 Preliminaries: Diffusion Models
Score-Based Diffusion Models [47, 48] learn a data distribution by coupling a noising (forward) Markov chain with a learned denoising (reverse) chain. Let $\{\bm{x}_{t}\}_{t=0}^{T}$ be the sequence of diffusion states with $\bm{x}_{0}$ drawn from the original data distribution $p_{\mathrm{data}}(\bm{x}_{0})$. In the forward process, Gaussian noise is added according to a fixed variance schedule $\bar{\alpha}_{t}$, where $\bar{\alpha}_{t+1}\leq\bar{\alpha}_{t}$, so that the learned marginal $q(\bm{x}^{T})$ approaches $\mathcal{N}(\bm{0},\bm{I})$ as $t\to T$. A score network $s_{\theta}(\bm{x}_{t},t)$ is trained to learn the score function $s_{\theta}(\bm{x}_{t},t)=\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{0})\approx\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t}|\bm{x}_{0})$, where the approximation holds under the assumption that $q(\bm{x}_{t}|\bm{x}_{0})$ is a close proxy for the true distribution $p(\bm{x}_{t}|\bm{x}_{0})$. The network is trained by applying noise $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$ perturbations to the $\bm{x}_{0}$, such that $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}},\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}},\epsilon,$ with the training objective minimizing the predicted error in the score estimate of the noisy samples:
@@ 3 Projected Langevin Dynamics
where the projection operator $\mathcal{P}_{\mathbf{C}}(\bm{x})=\operatorname*{argmin}_{\bm{y}\in\mathbf{C}}\|\bm{y}-\bm{x}\|_{2}^{2}$ returns the nearest feasible sample. Note that Langevin dynamic annealing occurs in an external loop decreasing the noise level across $t=T,\ldots,1$, so that noise level $\bm{\epsilon}$ and step size $\gamma_{t}$ remain fixed inside each subproblem. As $t\to 0$ the noise vanishes and the Langevin dynamics converge to a deterministic gradient ascent on the learned density function. Note also that annealed Langevin dynamics is known to reach an $\varepsilon$-stationary point of the objective with high probability [51]. In diffusion processes, the negative learned density function, $-\log q(\bm{x}_{t}|\bm{x}_{0})$, serves as the objective, so Equation (2) can be viewed as projected gradient descent on this landscape. The results of [51] extend to non-convex settings, giving further theoretical support to the constrained sampler adopted here.
@@ 4.2 Proximal Langevin Dynamics
$\mathbf{z}_{t}^{(i+1)}=\text{prox}_{\lambda\mathbf{g}}\underbrace{\left(\mathbf{z}_{t}^{(i)}+\gamma_{t}\nabla_{\mathbf{z}_{t}^{(i)}}\log q(\mathbf{z}_{t}|\mathbf{z}_{0})+\sqrt{2\gamma_{t}}\bm{\epsilon}\right)}_{\text{Langevin Dynamics Step}},$ (5)
@@
$\text{prox}_{{\lambda\mathbf{g}}}(\mathbf{z}_{t})=\arg\min_{\bm{y}}\,\bigl{\{}\mathbf{g}(\bm{y})+\tfrac{1}{2\lambda}\|\bm{y}-\mathbf{z}_{t}\|_{2}^{2}\bigr{\}}.$ (6)
@@ 4.3 Training-Free Correction Algorithm
Inner minimizer. Each proximal mapping is composed of a series of gradient updates to solve the proximal mapping. In practice, gradient descent is applied on the proximal operator’s objective:
@@
When a convergence criterion is met (e.g., $\mathbf{g}(\mathcal{D})(\mathbf{z}_{t}^{(i)})<\delta$), the algorithm proceeds to the next denoising step. Algorithm 1 provides a pseudo-code for this gradient-based approach to applying the proximal operator (lines 9-13) within the stable diffusion sampling process.
In the case that $\mathbf{g}$ is as an indicator function and a projection operator $\mathcal{P}_{\mathbf{C}}$ can be formulated in the image space, this can be simplified further be differentiating with respect to the objective of the projection, expressly $\|\mathcal{P}_{\mathbf{C}}(\mathcal{D}(\mathbf{z}_{t}))-\mathcal{D}(\mathbf{z}_{t})\|_{2}^{2}$. If $\mathbf{C}$ is convex, the projection can be constructed in closed-form (Section 6.1), but otherwise a Lagrangian relaxation can be employed [26]. For our experiments with non-convex constraint sets (Sections 6.2 and 6.3), we leverage an Augmented Lagrangian relaxation as described in Appendix D [19]. Notably, the gradients with respect to this projection’s objective function capture both the constraint violation term, $\mathbf{g}(\mathcal{D}(\mathbf{z}_{t}))$, and the distance term, $\tfrac{1}{2\lambda}\|\mathcal{D}(\bm{y})-\mathcal{D}(\mathbf{z_{t}})\|_{2}^{2}$, within the prescribed tolerance, resulting in the solution to Eq. (7) at the end of the minimization.
Importantly, in these cases the corrective step can be considered a projection of the latent under appropriate smoothness assumptions. We justify this claim with the following rationale: under the assumption of latent space smoothness
@@
[…]
for $t\leftarrow T$ to 00 do
@@
[24], the nearest feasible point in the image space closely corresponds to the nearest feasible point in the latent space. Therefore, while the distance component of the proximal operator, $\tfrac{1}{2\lambda}\|\mathcal{D}(\mathbf{z}_{t}^{i})-\mathcal{D}(\mathbf{z_{t}^{0}})\|_{2}^{2}$, is evaluated in the image space, minimizing this distance in the image space also minimizes the corresponding distance in the latent space, validating the use of this corrective step as a latent projector.
@@ 5 Complex Constraint Evaluation
While in the previous section we discuss how to endow mathematical properties within stable diffusion, many desirable properties cannot be directly expressed as explicit mathematical expressions. Particularly when dealing with physical simulators, heuristic-based analytics, and partial differential equations, it becomes often necessary to either (i) estimate these constraints with surrogate models or (ii) approximate the gradients of blackbox models. To this end, we propose two proxy constraint correction methods that leverage differentiable optimization to enforce constraints.
@@
Differentiating through blackbox simulators. In some cases (including the setting of our metamaterial design experiments in Section 6.2), a strong surrogate model cannot be leveraged to derive accurate violation functions. In such settings, the only viable option is to use a simulator to evaluate the constraints. However, these simulators are often non-differentiable, making it impossible to directly compute gradients for the proximal operator. To incorporate the non-differentiable simulator into the inner minimization process, this paper exploits a sensitivy analysis method inspired by the differentiable perturbed optimizer (DPO) adopted in the context of differentiable optimization [4, 33]. DPO operates by introducing controlled perturbations to the optimization variables and subsequently smoothing the objective function. We adapt this process by adding random local perturbations to the input parameters, evaluating the simulator’s output, and applying a smoothing function to approximate gradients. Then approximate gradients can be computed using a continuously differentiable Monte Carlo estimate,
$\mathbf{\bar{\phi}_{\epsilon}}(\bm{x}_{t})=\frac{1}{M}\sum^{M}_{m=1}\mathbf{\phi}\left(\bm{x}_{t}+\nu\epsilon^{(m)}\right)$
where $\phi$ is our external simulator, $\nu$ is a temperature parameter which scales the strength of the noise term $\epsilon$, $\mathbf{\phi}(\bm{x}_{t}+\nu\epsilon^{(m)})$ is a sample drawn from the Monte Carlo estimated distribution $\mathbf{\bar{\phi}_{\epsilon}}(\bm{x}_{t})$, and $M$ is the number of perturbed samples generated for the estimate. To obtain an appoximate gradient, the process differentiates with respect to this Monte Carlo estimate and formulates a loss:
$\nabla_{\bm{x}_{t}}\mathcal{L}\left(\mathbf{\phi}\left(\bm{x}_{t}\right)\right)=-\left(\bar{\mathbf{\phi}}_{\epsilon}\left(\bm{x}_{t}\right)-\textit{target}\right),$
that can be directly used to update the latent variable $\bm{x}_{t}$ through our proximal Langevin dynamics.
@@ 6 Experiments
The performance of our method is evaluated on three domain-specific tasks, highlighting its applicability to diverse settings. Supplementary results are provided in Appendix E
@@
    Projected Diffusion Models (PDM): We leverage a standard DDPM model denoising over the image space. Following [11], we project at each denoising step to restore feasibility. Importantly, this approach provides the current state-of-the-art for Section 6.1.
@@
    Conditional Diffusion Model (Cond): We apply a stable diffusion text-to-image framework, in order to condition the image generation on the constraint set.
When applicable (e.g., Section 6.2), we also augment our set of baselines with additional state-of-the-art models for the respective domains.
@@ 6.1 Microstructure Generation
Model MSE []delimited-[][$\downarrow$][ ↓ ] Fraction of physically invalid shapes []delimited-[][$\downarrow$][ ↓ ]
@@
Figure 4: Successive steps of DPO are shown. At each stage, $M$ perturbed shapes are generated, each undergoing structural analysis with $\phi$, which provides the corresponding stress-strain.
@@ 6.2 Metamaterial Inverse Design
Due to the complexity of the stress-strain response constraints in this problem, other constraint-aware methods (i.e. Projected Diffusion Models) are inapplicable, and, hence, our analysis focuses on the performance of Conditional Diffusion Model baselines. We compare to (1) an unconstrained stable diffusion model identical to the one used for our method and (2) state-of-the-art method proposed by Bastek & Kochmann [3], which operates in the ambient space. While our approach optimizes samples to arbitrary levels of precision, we observe that these baselines exhibit high error bounds relative to the target stress-strain curves that are unable to be further optimized. As shown in Figure 4, with five DPO steps our method provides a 4.6x improvement over the state-of-the-art model by [3] and a 5.1x improvement over the conditional stable diffusion model MSE between the predicted structure stress-strain response and target response. These results demonstrate the efficacy of our approach for inverse-design problems and generating samples that adhere to the target properties.

### v4

@@ Abstract
Stable diffusion models represent the state-of-the-art in data synthesis across diverse domains and hold transformative potential for applications in science and engineering, e.g., by facilitating the discovery of novel solutions and simulating systems that are computationally intractable to model explicitly. While there is increasing effort to incorporate physics-based constraints into generative models, existing techniques are either limited in their applicability to latent diffusion frameworks or lack the capability to strictly enforce domain-specific constraints. To address this limitation this paper proposes a novel integration of stable diffusion models with constrained optimization frameworks, enabling the generation of outputs satisfying stringent physical and functional requirements. The effectiveness of this approach is demonstrated through material design experiments requiring adherence to precise morphometric properties, challenging inverse design tasks involving the generation of materials inducing specific stress-strain responses, and copyright-constrained content generation tasks. All code has been released at https://github.com/RAISELab-atUVA/Constrained-Stable-Diffusion.
@@ 1 Introduction
However, despite their success in generating coherent content, diffusion models face a critical limitation when applied to domains that require outputs to adhere to strict criteria. In scientific and engineering contexts, generated data must go beyond merely resembling real-world examples; it must rigorously comply with predefined specifications, such as physical laws, safety standards, or design constraints grounded in first principles. When these criteria are not met, the outputs may become unreliable, unsuitable for practical use, or even hazardous, undermining trust in the model’s applicability. Bridging this gap is crucial for realizing the potential of diffusion models in scientific applications.
Recent research has reported varying success in augmenting model training with (often specialized classes of) constraints and providing adherence to desired properties in selected domains [12, 13, 14]. Many of these methods, however, are restricted to simple constraint sets or feasible regions that can be easily approximated, such as a simplex, L2-ball, or polytope. These assumptions break down in scientific and engineering tasks where constraints may be non-linear, non-convex, or even non-differentiable. Others are fundamentally limited, as training-time enforcement offers only distribution-level adherence to constraints rather than per-sample guarantees, even in convex settings, and cannot generalize to unseen or altered constraints without retraining. While inference-time enforcement methods address these shortcomings, these approaches modify the reverse processes in the original data space [15, 16]. This makes them incompatible with diffusion models like Stable Diffusion, which operate on learned lower-dimensional representations of the data. Latent-space variants have begun to appear, but they rely on special measurement operators [17, 18, 19] or on learned soft penalties [20, 21], thus limiting general applicability to the problems of interest in this work.
@@ 2 Preliminaries: Diffusion Models
Score-Based Diffusion Models [22, 23] learn a data distribution by coupling a noising (forward) Markov chain with a learned denoising (reverse) chain. Let $\{\bm{x}_{t}\}_{t=0}^{T}$ be the sequence of diffusion states with $\bm{x}_{0}$ drawn from the original data distribution $p_{\mathrm{data}}(\bm{x}_{0})$. In the forward process, Gaussian noise is added according to a fixed variance schedule $\bar{\alpha}_{t}$, where $\bar{\alpha}_{t+1}\leq\bar{\alpha}_{t}$, so that the learned marginal $q(\bm{x}^{T})$ approaches $\mathcal{N}(\bm{0},\bm{I})$ as $t\to T$. A score network $s_{\theta}(\bm{x}_{t},t)$ is trained to learn the score function $s_{\theta}(\bm{x}_{t},t)=\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{0})\approx\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t}|\bm{x}_{0})$, where the approximation holds under the assumption that $q(\bm{x}_{t}|\bm{x}_{0})$ is a close proxy for the true distribution $p(\bm{x}_{t}|\bm{x}_{0})$. The network is trained by applying noise $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$ perturbations to the $\bm{x}_{0}$, such that $\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ with the training objective minimizing the predicted error in the score estimate of the noisy samples:
@@ 3 Projected Langevin Dynamics
where the projection operator $\mathcal{P}_{\mathbf{C}}(\bm{x})=\operatorname*{argmin}_{\bm{y}\in\mathbf{C}}\|\bm{y}-\bm{x}\|_{2}^{2}$ returns the nearest feasible sample. Note that Langevin dynamic annealing occurs in an external loop decreasing the noise level across $t=T,\ldots,1$, so that noise level $\bm{\epsilon}$ and step size $\gamma_{t}$ remain fixed inside each subproblem. As $t\to 0$ the noise vanishes and the Langevin dynamics converge to a deterministic gradient ascent on the learned density function. Note also that annealed Langevin dynamics is known to reach an $\varepsilon$-stationary point of the objective with high probability [27]. In diffusion processes, the negative learned density function, $\log q(\bm{x}_{t}|\bm{x}_{0})$, serves as the objective, so Equation (2) can be viewed as projected gradient descent on this landscape. The results of [27] extend to non-convex settings, giving further theoretical support to the constrained sampler adopted here.
@@ 4.2 Proximal Langevin Dynamics
$\bm{x}_{t}^{(i+1)}=\text{prox}_{\lambda\mathbf{g}}\underbrace{\left(\bm{x}_{t}^{(i)}+\gamma_{t}\nabla_{\mathbf{z}_{t}^{(i)}}\log q(\bm{x}_{t}|\bm{x}_{0})+\sqrt{2\gamma_{t}}\bm{\epsilon}\right)}_{\text{Langevin Dynamics Step}},$ (5)
@@
$\text{prox}_{{\lambda\mathbf{g}}}(\bm{x}_{t})=\arg\min_{\bm{y}}\,\bigl\{\mathbf{g}(\bm{y})+\tfrac{1}{2\lambda}\|\bm{y}-\bm{x}_{t}\|_{2}^{2}\bigr\}.$ (6)
@@ 4.3 Training-Free Correction Algorithm
Inner minimizer. Each proximal mapping is composed of a series of gradient updates to solve the outer minimizer. In practice, gradient descent is applied on the proximal operator’s objective:
@@
When a convergence criterion is met (e.g., $\mathbf{g}(\mathcal{D})(\mathbf{z}_{t}^{(i)})<\delta$), the algorithm proceeds to the next denoising step. Algorithm 1 provides a pseudo-code for this gradient-based approach to applying the proximal operator within the stable diffusion sampling process.
@@
return $\text{violation}+\text{distance}$;
for $t\leftarrow T$ to 0 do
@@
In the case that $\mathbf{g}$ is an indicator function,
$\mathbf{g}(\bm{y})=\begin{cases}0,\qquad\bm{y}\in\mathbf{C},\\ \infty,\quad\;\;\bm{y}\notin\mathbf{C},\end{cases}$
the proximal mapping reduces to a projection onto the constraint set. If $\mathcal{P}_{\mathbf{C}}$ can be formulated in the ambient space, the minimization can be fully imposed on the projection objective, $\|\mathcal{P}_{\mathbf{C}}(\mathcal{D}(\mathbf{z}_{t}))-\mathcal{D}(\mathbf{z}_{t})\|_{2}^{2}$. Notably, the gradients of this objective capture both the constraint violation term, $\mathbf{g}(\mathcal{D}(\mathbf{z}_{t}))$, and the distance term, $\tfrac{1}{2\lambda}\|\mathcal{D}(\bm{y})-\mathcal{D}(\mathbf{z_{t}})\|_{2}^{2}$, within the prescribed tolerance, resulting in the solution to Eq. (7) at the end of the minimization. When $\mathbf{C}$ is convex, the projection can be constructed in closed-form (Section 6.1), but otherwise a Lagrangian relaxation can be employed [32]. For our experiments with non-convex constraint sets (Sections 6.2 and 6.3), we leverage an Augmented Lagrangian relaxation as described in Appendix D [33].
Importantly, in these cases the corrective step can be considered a projection of the latent under appropriate smoothness assumptions. We justify this claim with the following rationale: under the assumption of latent space smoothness [34], the nearest feasible point in the image space closely corresponds to the nearest feasible point in the latent space. Therefore, while the distance component of the proximal operator, $\tfrac{1}{2\lambda}\|\mathcal{D}(\mathbf{z}_{t}^{i})-\mathcal{D}(\mathbf{z_{t}^{0}})\|_{2}^{2}$, is evaluated in the image space, minimizing this distance in the image space also minimizes the corresponding distance in the latent space, validating the use of this corrective step as a latent projector.
@@ 5 Complex Constraint Evaluation
While in the previous section we discuss how to endow mathematical properties within stable diffusion, many desirable properties cannot be directly expressed as explicit mathematical expressions. Particularly when dealing with physical simulators, heuristic-based analytics, and partial differential equations, it becomes often necessary to either (i) estimate these constraints with surrogate models or (ii) approximate the gradients of black-box models. To this end, we propose two proxy constraint correction methods that leverage differentiable optimization to enforce constraints.
@@
Differentiating through black-box simulators. In some cases (including the setting of our metamaterial design experiments in Section 6.2), a strong surrogate model cannot be leveraged to derive accurate violation functions. In such settings, the only viable option is to use a simulator to evaluate the constraints. However, these simulators are often non-differentiable, making it impossible to directly compute gradients for the proximal operator. To incorporate the non-differentiable simulator into the inner minimization process, this paper exploits a sensitivity analysis method inspired by the differentiable perturbed optimizer (DPO) adopted in the context of differentiable optimization [35, 36]. DPO introduces controlled perturbations to the optimization variables and smooths the resulting objective, yielding differentiable surrogate gradients. Following this idea, random local perturbations are injected into the simulator inputs and define a smoothed function $\bar{\phi}_{\nu}(\bm{x}_{t})=\mathbb{E}_{\epsilon}\!\left[\phi(\bm{x}_{t}+\nu\epsilon)\right],$ where $\phi$ is the external simulator, $\epsilon\sim\mathcal{N}(0,I)$ is a random perturbation, and $\nu$ is a temperature parameter controlling the smoothing scale. An unbiased Monte Carlo estimate of $\bar{\phi}_{\epsilon}$ is obtained as
$\bar{\phi}_{\epsilon}(\bm{x}_{t})=\frac{1}{M}\sum_{m=1}^{M}\phi\!\left(\bm{x}_{t}+\nu\epsilon^{(m)}\right),$
where $M$ is the number of perturbed samples. The gradient of this smoothed objective can be written as
$\nabla_{\bm{x}_{t}}\bar{\phi}_{\nu}(\bm{x}_{t})=\frac{1}{\nu}\,\mathbb{E}[\phi(\bm{x}_{t}+\nu\epsilon)\,\epsilon],$
which corresponds to a finite-difference estimator whose scaling term $1/\nu$ is absorbed into the proximal step size during optimization. Finally, this smoothed estimator enables a differentiable loss formulation:
$\nabla_{\bm{x}_{t}}\mathcal{L}\!\left(\phi(\bm{x}_{t})\right)=-\left(\bar{\phi}_{\epsilon}(\bm{x}_{t})-\textit{target}\right),$
which is used to update the latent variable $\bm{z}_{t}$ (via $\bm{x}_{t}=\mathcal{D}(\bm{z}_{t})$) through proximal Langevin dynamics.
@@ 6 Experiments
The performance of our method is evaluated in three domains, highlighting its applicability to diverse settings. Supplementary results and baselines are discussed in Appendices E and F.
@@
    Conditional Diffusion Model (Cond): To assess the contribution of latent diffusion itself, we include a reference baseline consisting of an identical Stable Diffusion with text-conditioned constraints guidance [37].
@@
    Projected Diffusion Models (PDM): Following [15], this approach enforces feasibility by projecting onto the constraint set at each denoising step in the image space. PDM represents the current state-of-the-art for constrained generation in Sections 6.1 and 6.3.
  3. 3.
    Bastek and Kochmann (2023): For the task in Section 6.2, where PDM is inapplicable, we compare with this specialized method, which constitutes the state-of-the-art in that domain [38].
Collectively, these baselines capture the strongest existing constrained-generation methods, enabling a clear assessment of the improvements introduced by our latent constrained diffusion framework.
@@ 6.1 Microstructure Generation
Results. A sample of the results of our experiments is presented in Figure 2. Compared to the Projected Diffusion Model (PDM), latent diffusion approaches show a significant improvement. Latent diffusion models enable higher-quality and higher-resolution images. The previous state-of-the-art PDM, which operates without latent diffusion, had an FID more than twice as high as the models incorporating latent diffusion. Alternatively, the Conditional Diffusion Model, utilizing text-to-image conditioning, reports the best average FID of 10.8 but performs poorly with regard to other evaluation metrics. Conditioning via text prompts proved unsuitable for enforcing the porosity
Model MSE $\downarrow$] Fraction of physically invalid shapes $\downarrow$]
@@
Figure 4: Successive steps of DPO. The sample is iteratively improved and the stress-strain curve aligns with the target. Structural analysis shows progressive deformation under controlled compression.
@@ 6.2 Metamaterial Inverse Design
Due to the complexity of the stress-strain response constraints in this problem, other constraint-aware methods (i.e. Projected Diffusion Models) are inapplicable, and, hence, our analysis focuses on the performance of Conditional Diffusion Model baselines. We compare to (1) an unconstrained stable diffusion model identical to the one used for our method and (2) state-of-the-art method proposed by Bastek and Kochmann [38], which operates in the ambient space. While our approach optimizes samples to arbitrary levels of precision, we observe that these baselines exhibit high error bounds relative to the target stress-strain curves that are unable to be further optimized. As shown in Figure 4, with five DPO steps our method provides a 4.6x improvement over the state-of-the-art model by [38] and a 5.1x improvement over the conditional stable diffusion model MSE between the predicted structure stress-strain response and target response. These results demonstrate the efficacy of our approach for inverse-design problems and generating samples that adhere to the target properties.
6.3 Copyright-Safe Generation
Results. Figure 5 (bottom-right) reports the FID and constraint satisfaction, and Figure 5 (top-left) compares the evolution of the original sample and corrected sample. We implement a Conditional Diffusion Model baselines using and unconstrained stable diffusion model identical to the one used for our method. The conditional baseline generates the protected cartoon character (Mickey Mouse) 33% of the time, despite conditioning it against these generations. The Projected Diffusion Model also struggles in this domain, particularly due to the image-space architecture’s inability to handle high dimensional data as effectively as the latent models; it failed to generate reasonable samples at a resolution of $\times1024$, unlike the latent models which maintained image quality at this scale. As a result, we were constrained to operate at a much lower resolution of $\times64$ during sampling, subsequently relying on post-hoc upscaling techniques to reach the target resolution. This led to higher FID scores and only marginal improvement over the conditional model’s constraint satisfaction. Conversely, our Latent Constrained Model only generates the protected cartoon character 10% of the time, aligning with the expected bounds of the classifier’s predictive accuracy. Our method has proven to be highly effective because it preserves the generative capabilities of the model while imposing the defined constraints. The FID scores of the generated images remain largely unaltered by the gradient-based correction. This demonstrates that our approach can selectively modify generated content to avoid copyrighted material without compromising overall image quality.
7 Conclusion